Constants Part A A mass is oscillating with amplitude A at the end of a spring How far (in terms of A) is this mass from the equilibrium position of the spring when the elastic potential energy equals the kinetic energy? d- Submit Request Answer Return to Assignment Provide Feedback
Answer #1

Part A:

We know,

The potential energy in a spring,P.E = (1/2) kx2

The kinetic energy in a spring, K.E = (1/2) mv2

he total energy of syetem = P.E + K.E

E =(1/2) kx2 + (1/2) kx2. (because, if P.E = K.E)

E = kx2

Total potential energy occurs when x = A.   

E = kA2 = d